In a 70g 15 percent solution how much of a 90 percent solution does it take to make an 80 percent solution?
Please show me how to set up the
Problem and solve it too. Thanks

Answer by stanbon(75887)

In a 70g 15 percent solution how much of a 90 percent solution does it take to make an 80 percent solution?
------
Equation:
active + active = active
0.15*70 + 0.90x = 0.80(70+x)
------
Multiply thru by 100 to get:
15*70 + 90x = 80*70 + 80x
10x = 65*70
---
x = 7*65
x = 455 grams (amt. of 90% solution needed)
---------------
Cheers,
Stan H.